OpenStudy (anonymous):

a 12ft ladder is leaning against a wall.if the top of the ladder slips down at a rate of 3ft/s,how fast will the foot be moving from the wall when the top is 11ft above the ground?

OpenStudy (anonymous):

x² + y² = 12² -> x = .... Take derivative: => x' = - y y' / x

OpenStudy (anonymous):

okay i got that x equals the sqrt of (12^2 -y^2) now the derivative of of that would be -y(2y)/sqrt(12^2-y^2)???

OpenStudy (anonymous):

nope i did something wrong.....

OpenStudy (anonymous):

x = √ (12² - 11²) = ...

OpenStudy (anonymous):

okay but then thats a numbeer so the derivative of it would be zero?....

OpenStudy (anonymous):

=> x' = - y y' / x NOW, it's time to plug in

OpenStudy (anonymous):

so the answer is -50.46?

OpenStudy (anonymous):

no thats impossible what am i doing wrong?

OpenStudy (anonymous):

I'm getting so tedious of you showing nothing instead of random number, then "..wrong"

OpenStudy (anonymous):

okay well i took x'=-y y'/x and put in 11 for y and then the derivative for y would be 2y so its 22 so so far i have (-11)(22)/x and then for x i plugged in sqrt(12^2-11^2) so i had (-11)(22)/sqrt(12^2-11^2) and from that i got -50.46.......

OpenStudy (anonymous):

Can you even find this x = √ (12² - 11²) = ... x' = -y * y'/x = - 11 ( -3) / ...

OpenStudy (anonymous):

1.4347

OpenStudy (anonymous):

12² = ... 11² = ... 12² - 11² = ... => √ (12² - 11²) = ...

OpenStudy (anonymous):

You have serious problem with basic arithmetic !!!

OpenStudy (anonymous):

12^2=144 11^2=121 144-121=23 sqrt(23)=4.7958

OpenStudy (anonymous):

See, at least you can punch some numbers into your calculator (sighs )

OpenStudy (anonymous):

x' = - 11 ( -3) / √23

OpenStudy (anonymous):

im on here to get help not critized obviously im not good at this thats why im asking the questions not helping people....6.881

OpenStudy (anonymous):

Look I already done everything, and I don't understand how hard it is just to put in the numbers into your calculator !!!

OpenStudy (anonymous):

okay first of all its i aldready did everything not done...and im sorry it doesnt make sense and you didnt already do everything you gave me a bunch of letters that mean nothing to me and with no paranthesis